OpenStudy (anonymous):

a 1.00 kg toy car travels at a constant speed in a uniform circle on flat horizontal surface. The car completes one circle every 3 seconds and the radius of the circle is 2 m. What is the magnitude of the net force that acts on the car? (a) zero (b)8.77N (c) 4.19 N s (d)0.222 N (e) there is no sufficient information

OpenStudy (perl):

we can use Fnet = mv^2 /r , also since v = w*r we have Fnet = m* (w*r)^2/r = m * w^2 * r

OpenStudy (perl):

so w = 2*pi rad/ 3 sec

OpenStudy (perl):

Fnet = (1 kg ) * (2*Pi / 3 ) ^2 * 2 = 8.77 Newtons

OpenStudy (anonymous):

oohh makes since now i didnt know to solve for v thank you!!